Restart Night #14 - UT 04 Sep 2020

OSA: J. Williams

ISp: E. Solheid

Mtn Manager: J. Schuh

SO/LBTI: J. Power

LBTI: S. Ertel

AO: G. Taylor, B. Rothberg (scripting and targets)

NOTE: Split Night between LBTI Restart and AO-SOUL-LUCI - Handover 00:14


Summary:

Completed basic onsky checkout for LBTI with a few hitches. Distortions noted int eh LMIRCam images. Played with NCPA which improved thigns but this may be easier closed dome. M3/M2 alignment determined for LBTI to align the pupils. SOUL AO checkout started. closed on bin1 and bin4 targets. Completed a science script with large throws without issue. And obtained OSCO data.

Problems:

LBTI: Variable seeing and distortion in the LMIRCAM images made it difficult to optimizer the focus between the WFS and LMIRCAM. Played with NCPA. Ok but not optimal yet. From recently installed pupil mirror?

Proposed Plan:

  • First half LBTI
    • (DONE bint 1 only)Close loop on bin 1 and bin4
    • DONEverify tertiary postion and pupil alignment
    • DONErun a focus sequence to verify the focus on LMIRCam
    • run through a quick science script to verify communication
  • If LBTI finishes early:LBC_checkout_targets
    • LBC-Bino Checkout
      • rotator center measurements for both LBCs:
        • Doing this at 2-3 different elevations would be ideal.
        • Slew to a copointing field and collimate and copoint
          • Any copointing script can be used. A jskycalc-formatted catalog of coords for the Sci-like OBs is in Restart_2020/LBC/LBC_checkout_targets.cat and a visibility plot is LBC_checkout_targets_vis.png
        • Run the series of OBs: ~/Service_Observing/Restart_2020/LBC/RotCen/pa0.ob, pa180.ob, pa90.ob and pa270.
          • These OBs (paxx.ob) have been hacked to set DEC to -90 so that a new preset is not sent and the instrument is only rotated.
      • collimation
        • super-focus sequence for both
        • superfoc at 2-3 different elevations would be ideal.
  • SOUL/LUCI/AO 2nd half
    • System Checkout
      • Diffraction-Limited Verification
        • DONEVerify Bin 1 and Bin 4 (2hrs)
        • DONETest Science Script M52 (1-1.5 hrs)
        • DONEOSCO (0.5 hr)
      • ESM verification
        • Test Science Script M52 (1.5 hr)
      • Poor Weather Contingency
        • ESM Verification (as long as loops can close collect data)
        • Check WFS magntiudes against known magnitudes

Nightlog:

1:10 Doug has verified the LUCI transforms. Copied PCSInstrument.conf.20200902_2 with the updated RIGHTZEROPOINT of 176.95 (updated from 177.05). Josh has restarted PCS.

1:15 Bringing up LBTI WFS. LBTI Log generated:
https://docs.google.com/spreadsheets/d/1hY7nBicp66ONE7qWn-ezrZibhhv_wUO-Sucv-TgNbyU/edit?usp=sharing

2:08 Pointing and closing on Vega. First reoptimization looked not great. pupil alignment not ideal and there looks like there is some slow seeing tonight.

2:22 aligning M3/M2 on LMIRCAM
  • Absolute positions to start tip/tilt +200/+300
  • Final Absolute -115/+250 (Global Offsets -315, -50)
2:41 Ajusting bayside z focus for LMIRCAM
  • running bayside z from 79 - 72. Optimal focus looks to be around 76. Going to execute a more narrowed focus sequence around z=76. We are at 83 in elevation and vega is very bright, hard to get a non saturated image. Moving to somethign fainter for refined focus.
2:51 Moving to 11Acq

Commanded optimize gain but adsec arbitrator did not receive? Restarted the WFS Arb, opened loop and closed manually to kick things back into talking and synced back up. All happy again.

Stepping from z=76 in steps of -0.1, focus appears to be 75.3 but not clear cut. Reflattening and reoptimizing. there appears to be somethign at 8:30 on the OCam. Not showing on mirror GUI, somethign vignetting? rehomes the fw's, baysides,

We are struggling to find optimal focus on LMIRCam. Homing the ADC's. ADC2 controller activing weird and unstable. But ADC positions look more like what we expect for current telescope position after ADC controllers restarted and rehomed. ADC's did not make a difference

stepping from 73.5-77.5 and determining optimal focus. Airy rings of long integrations continue to look bad/distorted even at the "best" focus.
Taking way out of focus z=55. Looks like expected with some trefoil..
Running through with long exposures 75.6 gives the best shape but distored Airy rings, possibly the result of the new pupil mirror? Core looks good. Perhaps we will try some NCPA. Playing with NCPA
Best values are still not very good. The PSF has some shape
Z5=-9 , Z6=12 Z7=14 Z8=-4 Z10=-8 Z11=7 Z12=6 Z22=-4, z=75.6
Focus, tip tilt? We should try to focus with the argos source?
Testing dithering script with LMIRCAM. Failed. Script failed to wait for AO before taking LMIRCAM data.
Bayside Z focus in the LBTI_ACE-AO in /home/sxwunit/aoroot/calib/wfs/current/setups to 75.6

7:17 Handing over for SOUL AO

7:23 Slewing to first OSCO target: HD216399
  • I left the FLAO gui in intervention mode when i finished up LBTI...took us a minute to catch that...
  • Closed perfectly and the graphs even showed in FLAO.
  • luci1.20200904-0001-00020
7:39 Greg is sending presets to nearby bin1 and bin4 verification targets.
  • AO32 HD1480 Bin1 verification
  • When applying optical grain correction the star took a little walk. Troubles clearing the preset, then the pointing was off when resent.
  • Scripted teh bin1 data gathering
  • luci1.202020904..33 dark
  • 0034-0038 FeII/ND2 1x10x2.5 Cube (-5,5)
  • 0039-0043 FeII/ND2 1x10x2.5 Cube (-5,5)
data looks good, bin1 gets Greg's seal of approval: scrumptous

8:45 Sending bin4 verification taget, AO91. LHS1443
  • target hit bin3
  • luci1.20200904.0044-0053 10x3x2.51 blind blind
  • ADSEC rip. Greg ended up having to adsec stop start. Reason for RIP not really known, just a cranky adsec when opened loop.
9:08 bin4 AO115 is a little fainter, going now. POinting was off when we got there so Joch is doing a pointing check.
  • Struggling to find a bin4 star.
  • StoneA1-17, 16.99 according to stoneA catalog but still went to bin3.
  • General seach around, N32013214476 R=17.35.
  • Background from moon keeps kicking us to bin3 (atleast thats the guess). Moving north.
Slewing to GSC Catalog star with Rmag of 17.47. Estimated magnitude on OCAM is 18.
  • The camera lens hit its limit, moving the rotator PA+90
  • We rock, bin 3 with 16.3, and bin 4 18.0. Not too shabby. "Booyah"
  • The luci image is a little elongated, bit looked decent.
  • Luci1.20200904.0053-0058 bin4 on R17.35 (18.0 on OCAM, 20 average pupil intensity ~20, 3.8 photons /subapp)
10:59 Moving to M52.
  • Pointing check needed, also aligning field stop. executing K only. Bin2, freq 849, 250 modes. Loop opened on very first image (0060). Greg is reoptimizing.
  • Failed on first offset, likely because greg reclosed manually. Resending preset.
  • first image luci1.20200904.0061
  • 0061-0085 K images on M52. No issues with anything once we were up and running
12:15 Closing

Obtaining darks and N30 flats
  • Camera failed to reach position N375 when script commanded. fixed
  • Fieldstop misaligned 0126-0135
    N30Flats: J, H, K, Ks, FeII 0140-0190


-- BarryRothberg - 03 Sep 2020
Topic revision: r13 - 04 Sep 2020, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback